Oxford Materials IT Committee Meeting

MICHAELMAS TERM 2016

There will be a meeting at 11am on Tuesday 25th October 2016 in Banbury Road Committee Room..

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting:  Tuesday 31st January 2017 at 11am in BR Committee Room 

ITEM 2. MATTERS OUTSTANDING FROM THE PREVIOUS MEETING

See minutes below.

ITEM 3: MATTERS FOR DISCUSSION

Update on recent infrastructure outages (attached document)

Verbal reports on the status of various projects:

IT services projects: TONE, Chorus, OpenAccess
Departmental projects: Server infrastructure, Photocopiers, GreenImpact , iLab
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group

ITEM 4: HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Re-advertising the Grade 6 IT staff position is still on hold pending the outcome of a wider Review of IT. After numerous delays, IT services carried out the review in June/July however the reviewer changed job in August and the final report has not been forthcoming! 

Admin Network Lots of staffing changes over the summer and associated changes in accounts/access rights/computers. New Head of Administration and Finance Charlotte Sweeney started in August – taking over existing desktop but purchased new laptop. Barry Fellows retired and was replaced by Debbie Townsend. Rosalind Sainty is temp cover for the vacant finance assistant position. Karen Disney retired and was replaced by Rebecca Bradford. Jane Foxall has taken over Rebecca Bradford’s role. Louise Hakansson is temp cover for the vacant receptionist position. Marije Zeldenrijk left and has been replaced by temp Lorraine Maggi. Ian Sutton now spends time working in the finance office as well as being storesman.

Teaching Laboratory All computers were checked/patched before the start of the new academic year. The TL colour printer was decommissioned and removed, moving all undergraduate printing onto the nearby photocopier. PJW demonstrated the first year practical for the 36 new students.

Digital Signage: Electronic display signage continues to be updated by Lorraine Laird. Over the summer the Begbroke display signs went offline due to power rewiring works. The device has been mended but still needs returning to Begbroke.

Lecture theatres / meeting rooms AV facilities were majorly renewed/upgraded in HRLT and ETBCR, implementing HDMI and WiPG wireless presentation. Also Skype functionality has been implemented in ETBCR and BRCR. The BRLT upgrade was postponed until next financial year due to budget constraints. Still need to fit new roller whiteboards in HRLT and ETBCR but dependent upon maintenance staff.

Printers and photocopiers: The TL printer was decommissioned, however AOT blocked decommission of the admin colour printer. Ricoh have been extremely unhelpful processing the annual papercut license renewal which is still outstanding from July.

Digital Print-room IT staff have purchased three years’ worth of hardback thesis binding covers in order to take advantage of a “free binding machine” offer, thus replacing the unit which failed a couple of years ago and was not replaced due to cost. (binding machine also in reception).

Library The library2 computer is in need of replacement with trickledown hardware.

MML Four replacement motherboards were purchased after the January aircon failure. Ironman continues to have issues. The aircon 50% failed irreparably in April. PJW obtained quotes for complete replacement of all aircon for £16k+VAT however the department has been unable to fund the work. Over the summer the remaining 50% aircon tripped off on ~5 occasions and another motherboard was damaged

FIM International workshop was held at StCatz in September using posterboards. 

EM Some space was cleared from the EM storage array. AJW is purchasing further capacity. Gareth Hughes has been testing new data analysis system. ARM was upgraded with additional computers. Several EMs were decommissioned from Begbroke or relocated to central site.

MFFP Monthly patching and rebooting of 13 systems and storage arrays. MFFP website is desperate for upgrading due to Drupal6 end of life last April. Old disks continue to fail and be replaced periodically. The storage disaster failover needs testing again.

OMCS Three new UPS systems were commissioned (Northstar/EPMA/INCA)

PGB Salto door entry system still needs fixing to work with university cards rather than separate cards. Steady stream of new people and equipment.

Website Daily/weekly editing of department website and news items. New 64bit webserver (second-hand hardware) was commissioned to replace end of life old 32bit webserver. Over the summer 20 sites were migrated, but more than 30 still need migrating, mostly due to CmsMadeSimple needing upgrading 1.13 to 2.1 and configuring for https and where possible incorporating responsive design. MFFP website needs migrating to Drupal7.

Publications Weekly editing of Departmental Newsletter. 9 sets of business cards. Signage for opendays (July/Sept) and workshops. Posterboards for workshops (Castep, Atomprobe, NQIT, QIP, Dept60th). Printed handbooks for PartII, Prelims, Graduate. Still awaiting Finals and Dept handbooks.

Network As part of the university TONE and Odin project, feeding a new fibre into Hume-Rothery building was not possible due to the discovery of asbestos on the main (and only) route to the data cabinet. Unfortunately, this means that no new network connections can be made within the Hume-Rothery building (or until a new route is created). As a workaround some unused fibres in the existing cable were terminated and the department migrated to the new university backbone on 8th May. All other departmental buildings have also had their FroDo units migrated to the new backbone. New wiring was installed from the newly installed FroDo cabinet in BR to feed the other cabinets thus allowing the wireless access points in BR to be upgraded and additional access points added. New network wiring was installed in ETB level10 for Dave Armstrong’s new lab and level 30 for Lapo Bogani’s new labs.

Security Oxcert released 55 security bulletins. Several minor virus infections dis-infected. The university continues to be targeted with phishing emails which many people are reporting. (no-one has reported falling for phishing). A new departmental windows update server was implemented in order to provide support for Windows10 updates. The new server is currently managing 60 systems so far (including 10 win10 systems) and IT staff are gradually migrating the >500 other systems and resolving problems as they arise

WDE Still only six user’ devices encrypted – need to expand uptake targeting senior staff. Limited by time/manpower…..

Disposal In May a WEEE skip was used to clear Richard Cripps’ office and dispose of other miscellaneous IT equipment.  In October the WEEE “postoffice” storeroom was emptied (due to reallocation to Clara Barker)  filling a van, including disposing of ~20 CRT/TVs and onsite crushing of 178 hard disks.

Software The Origin university site license managed by PJW was renewed for 3 years and Origin2016 is being deployed. New version of Abaqus2016 is being deployed (significant changes). Windows10 anniversary update was released in August and is a recommended upgrade. IT staff are still deploying a mixture of Windows7 and Windows10 systems. MacOS Sierra 10.12 was released in September but is still NOT recommended due to third party compatibility issues (Office, TSM) but is becoming unavoidable for new purchases.

Disaster Recovery and Backup Several disk fails have occurred and data has been recovered or restored from backups, however several users have lost personal data due to not having backups. Aircon in the Hume-Rothery server room failed 50% in September. Funding is being sought to entirely replace the failed unit (possible central funding). PJW is still looking out for a second-hand replacement Symmetra UPS power module (<£1200) – currently this UPS is operating below capacity but is an acceptable risk. PJW still needs to finish rolling out purchased research group UPS’s into production (KP1/KP2/KP3/KAQOR).

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New people & accounts May 5; June 9; July 8; August 2; Sept 9; Oct 86;

New Device registrations May 6; June 16; July 8; August 7; Sept 18; Oct 38;


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 25th October 2016 in Banbury Road Committee Room.

The following people attended: CS, PJW, RS, CA, RT, SLP, Flynn Castles, Andy Naylor, Lewys Jones

1. Apologies for Absence and Membership of the Committee

Paul Bagot and Jiale Wang sent their apologies.

Paul Warren welcomed Charlotte Sweeney who joined the department as Head of Administration and Finance, thus taking over from Alana Davies as chair of the IT Committee. Committee membership on the website and committee mail list had been updated accordingly.

Paul Bagot will leave in November therefore the committee was asked to advise upon possible candidates to be appointed by HoD. Consideration of building representation focussed upon Parks Road therefore Jamie Warner or Lapo Bogani or Jan Mol were considered possible candidates. PJW will liaise with HoD.

Action: PJW/PSG

  2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *Draft Information Security Policy (Minute 2a)

Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy, supported by PSG.
There has been steady progress towards getting staff to complete the awareness module. In September the status was that of the academic and admin staff, 70 have completed the module whereas 22 are still outstanding even after monthly reminders.

Action: PJW/PSG

b) *Office upgrade (Minute 2b)

The admin Novell desktops use Office2010, originally due to compatibility with central admin macros. APD asked whether Office2013 or Office2016 could be rolled out to admin desktops yet. IT staff currently deploy Office2016 on non-admin desktops, however in August 2017 the whole university will be migrating to Office365 which is taking over from Nexus for provision of email, therefore the request to upgrade the admin desktops has been postponed to August 2017.

Action: PJW

c) *Infrastructure outage risks (Minute 3a)

After the outage in May 2016, APD requested that IT staff put together a document detailing the main areas of risk associated with single points of failure within departmental IT systems, and presenting costed proposals on how to minimize future risks. This document could then feed into the forthcoming Review of IT by IT services.

The requested document was not produced and the IT review has happened (but not reported). IT staff have attended a VMware training course (PJW/RS) and replacement switching for the VM infrastructure is being implemented to address performance issues. It is still desirable to have a document ready for consideration for budget planning in March.

Action: PJW/RS/AMC/CA

d) *Photocopier guidance (Minute 3b)

PJW wrote a Newsletter Mildred article (2016-05-20) about the photocopiers, including reference to the importance of solving/reporting manual feed paper jams in order to avoid unscheduled job release with the potential for sensitive data being viewed.

e) Print cost reporting  (Minute 3b)

In Hilary Term there was a recommendation to implement monthly email reports of print/copy usage to all users showing the costs in order to reduce usage.

The solution appears to be to email users to prompt them to login to the web interface at https://print.materials.ox.ac.uk:9192/user (since the Papercut software cannot automatically email usage reports for each user to each user, as opposed to emailing a report to a specific person). This can be done via a (termly?) newsletter item.

Action: PJW/LL

The motivation for this request was also discussed. LJ noted that comparative data was more useful that reports of absolute usage. Maybe summarise by building/group? CS queried the cost effectiveness of chasing this – people will print/copy what they need to - the motivation should be to promote greener behaviour but not to expend effort chase pennies.

f) *Teaching Lab printer  (Minute 3b)

The teaching laboratory colour printer was decommissioned and all TL computers were configured to print to the photocopier. Student feedback has been positive and there appears to be much less paper wastage since uncollected print jobs automatically deleted if not released.

g) *Finances (Minute 5)

PJW reviewed expenditure and submitted all stock journal requests prior to the financial year end on 31st July.

h) Display Signage (Minute 6a)

SLP requested that the departmental display signage system be extended into the Holder Building cafeteria in order to counter-balance the Engineering display signage.  IT staff have considered this request and identified a possible location above the chiller cabinet, however this expenditure (~£500) was not budgeted for this academic year and funds are tight..

Action: PJW

i) Begbroke Map  (Minute 6b)

RT noted that the visitor information leaflet needs some Begbroke information updating, including the building site map.  PJW still needs to do this..

Action: PJW

3 Matters for Discussion

PJW updated the committee verbally upon the status of the following.

IT services projects:

  • TONE – replacement backbone network. http://projects.it.ox.ac.uk/tone/ Project has been completed as far as this department is concerned. A few minor tasks are outstanding (removing old hardware in BR and faster link in 14PR).
  • Chorus – the new telephone system http://projects.it.ox.ac.uk/icp/ Also see official Chorus newsletters http://projects.it.ox.ac.uk/icp/newsletter All handsets in Materials will be replaced/migrated between April-June 2017. (now to be paid centrally – which helps Departmental budget). PJW is managing the pre-migration data collection and validation for each telephone line within the department (handsets/phonenumbers/people/directoryentries/networkports/powerplugs) and completing the network readiness questionnaire for each of our 13 buildings. The actual changeover is intended to be short and sharp (a team of 6 people are working around the university building by building). The enhanced software features of the new Chorus system, although mostly available, are not the target of the current migration project.
  • OpenAccess http://openaccess.ox.ac.uk/ This week is “OpenAccess Week” aiming to raise awareness that all publications need to be submitted via the Act of Acceptance process. PSG and PJW are the departmental OpenAccess monitors but systems are such that it is not easy to check that papers are being submitted properly. SLP noted that it would be useful to have a departmental web/weblearn page summarising the OpenAccess guidance notes that have been circulated via email.

    Action: PJW

  • Information Security Improvement Programme (ISIP) This department still needs to review and adopt an Information Security Policy.(Minute 2a). The HoD is likely to receive an Information Security Baseline Questionnaire that needs to be completed by January. The aim of the questionnaire is to gather status information rather than to try to fix everything immediately.

    Action: PJW/PSG

  • Replay (automatic lecture recording service) http://help.it.ox.ac.uk/replay/index The Replay service has ended pilot and entered production. At present it is still free. Several departments now routinely record all their lectures. This department could follow that route but would need video (to cover the chalk&talk) in addition to the baseline audio+powerpoint. PJW will discuss this with AOT for consideration by DMAC prior to budgeting for suitable video cameras for HRLT and BRLT. LJ also questioned whether this might make remote attendance to colloquia possible for people based at Begbroke and elsewhere.

    Action: PJW/AOT

  • Nexus Upgrade Office365 IT services are moving forward with the plan to replace Nexus Email service with Office365 online service (and 50Gb quotas) due to start in September 2017.

Departmental projects:

  • Server infrastructure – replacement switching has been ordered for the VMware infrastructure to try to reduce intermittent slowdowns associated with high IO. New webserver and new windows update servers have been commissioned.
  • Photocopiers – Printed to the photocopiers has been well received and is in widespread use. CS’s copier continues to behave erratically and needs attention. Next year’s budget should contain a line for adding a copier to 14 Parks Road.
  • iLab – CS has taken over from APD as primary contact with iLab. Three CFAS equipment have been added to the EM core. The Teaching Lab XRD which is due to move into a restricted room will probably not be added to the system too since that is not a Small Research Facility.
  • Green Impact – On 15 June 2016 The University of Oxford held the annual local Green Impact Awards at the Blavatnik School of Governance. The Department of Materials was awarded 'Gold' for general environmental practices, 'Silver' for the labs and a special 'Innovation' award for Milly, the mascot. PJW asked ITC members to promote the green impact activities or even consider joining the team.

Information from other committees:

  • MPLS ICT panel – the Michaelmas Term meeting yesterday received a status report from IT Services. Items of interest have all been covered above.
  • Divisional Software Licensing Group http://help.it.ox.ac.uk/sls/background Origin site licence was renewed in July for 3 years (managed by PJW). Matlab, Novell site licenses were also renewed. Nothing else to report here.

Having discussed the various IT Services projects, PJW did a quick straw poll among the committee to see whether people were aware the university operate a weekly “at-risk” period 7-9am each Tuesday morning for maintenance of IT systems. Only IT staff seemed aware of this, therefore PJW should have an annual Newsletter item.

Action: PJW/LL

4. Head of IT Report
A written report was circulated with the agenda.

RT noted the new wireless and requested an access point in BB Annex. PJW agreed to relocate a spare access point (removed from BR).

Action: PJW/RS/RT

5. Finances
No status report was presented. The budgeted lecture theatre upgrade £10k has been spent. IT staff are currently rolling out new desktops and laptops to the new students.

Action: PJW

 6. AOB

The process for replacement of photocopier toners via reception seems unreliable. PJW to liaise with Ricoh and push for more spares to be kept onsite rather than ordering when running low.

Action: PJW

7. Date of Next Meeting

The next meeting is Tuesday 31st January 2017 at 11am in 21 Banbury Road Committee Room.

The meeting closed at 12.20pm.